OpenStudy (anonymous):

evaluate the series 4-2+1-0.5+0.25 to s10. Round to the nearest tenth. a) 5.50 b)-2.75 c)2.75 d) 2.66 PLEASE HELP!

OpenStudy (anonymous):

just add up the first ten terms yes?

OpenStudy (anonymous):

i got 2.66

OpenStudy (anonymous):

\[4-2+1-\frac{1}{2}+\frac{1}{4}-\frac{1}{8}+...-\frac{1}{128}=2.664...\]